OpenStudy (anonymous):

Find all values of t in the interval [0, 2 π] that satisfy cos^2 t-4cost+3=0 . (Hint: Substitute x = cos t, factor the result, find the values of x that solve the equation, then put cos t back in for x.)

OpenStudy (anonymous):

Comment: The equation should be cos2 t - 4cos t + 3 = 0. Rewrite this as x2 - 4x + 3 = 0 and factor to solve for x. Now going backward cos t = the numbers you got for x. You should be able to find the correct answers from the Unit Circle.

OpenStudy (goformit100):

Options

OpenStudy (anonymous):

This one is short answer. I solved x2 - 4x + 3 = 0and got x=1 and x=3